OpenStudy (anonymous):

Donna has a new job. Her annual starting salary is $17,600. She will receive a raise of $850 each year. Which expression models her salary n years from now? What will Donna's salary be 5 years from now?

OpenStudy (thread252):

do you have an answer choice?

OpenStudy (anonymous):

Y=850x + 17600

OpenStudy (anonymous):

what is all your choise

OpenStudy (thread252):

year 1) 17600 year 2) 17600+850=18450 year 3) 18450+850=19300 year 4) 19300+850=20150 year 4) 20150+850=21000 so by the end of 5 years her annual pay check would be 21000 dollar.(that is a very small paycheck, what kind of job do she have)

OpenStudy (anonymous):

in my country that amount is big